# C-3DPO: Constrained Controlled Classification for Direct Preference Optimization

### v1

@@ 7 Related Work
A few recent papers have presented a unifying perspective on DPO-style algorithms. Notably, Tang et al. (2024) presented a generalization of DPO where different supervised learning losses are applied to the difference of implicit rewards, $\big{(}\hat{r}(x,y_{w})-\hat{r}(x,y_{l})\big{)}/\beta$. In contrast, we make a clear connection between DPO-style algorithms and classification, and also extend our results to lists and auxiliary information. A second notable example was to generalize from KL to any f-divergence when measuring the discrepancy between the target and reference model (Han et al., 2024). We also note that the first ingredient of our classification framework, namely $\pi_{\theta}$, was used by Sharifnassab et al. (2024) to propose a soft version of DPO.
As mentioned in Section 4, earlier works studied the undesirable decrease of the winner probability during training. These works explained this phenomenon from different angles, and thus, proposed different losses to address it. Here we provide a brief overview of a number of these results, especially those that we experimentally compare against. Pal et al. (2024) proposed DPOP which addresses the phenomenon by adding the penalty term $\max\big{(}0,-\hat{r}(x,y_{w})/\beta\big{)}$ within the log-sigmoid of the DPO loss (5). DPOP can also be viewed as DPO with a modified BT model. In this sense, it has similarities with the $\alpha$-DPO loss (Shen et al., 2024).
Xiao et al. (2024) attributed the undesirable behavior to the contrastive loss of DPO not being scale-calibrated, i.e., ignoring the absolute values of implicit rewards, $\hat{r}(x,y_{w})$ and $\hat{r}(x,y_{l})$. They address this by constraining the implicit rewards to a scale that matches the ground-truth rewards $r(x,y)$. Thus, in their proposed loss, Cal-DPO, they add the square-loss $\big{(}\hat{r}(x,y)-r(x,y)\big{)}^{2},\;y\in\{y_{w},y_{l}\}$ to the DPO loss (without $\beta$). Of course, since they do not have access to the ground-truth reward, they replace it with $1/2$ and $-1/2$ for $y_{w}$ and $y_{l}$, respectively. A loss very similar to Cal-DPO was proposed by Wu et al. (2024) and they named it SPPO. It is simply Cal-DPO without the DPO loss. Finally, D’Oosterlinck et al. (2024) proposed APO, which offer fine-grained control over the implicit rewards.
@@ 8 Conclusion & Future Work
In this work, we made two contributions to the preference optimization literature. First, we showed that many DPO-style algorithms can be viewed as classification algorithms. In this classification framework, one may choose a specific classification label (soft or hard, using auxiliary information or not) and a specific loss to obtain a DPO-style algorithm. We believe that revealing this connection promotes a mutual transfer of ideas between the classification and preference optimization.
The classification framework also revealed that the underlying problem solved in DPO-style algorithms is under-specified and therefore susceptible to the collapse of both winner-loser probabilities. By endowing the DPO-style algorithms with our newly proposed constraints, we showed that the problem becomes specified, and thus, probability-collapse can be avoided.
A curious observation we made is that while we do want to avoid probability collapse, it is also not true that increasing the probability of the winner will always lead to the best final model. This means that the best performance is usually achieved by increasing the probability of some of the unseen answers. Thus, in future, it is interesting to do a more systematic study of the kinds of unseen responses for which DPO-style algorithms increase the probability during training.
Appendix A Derivations of IPO as a Classification Algorithm
@@
Appendix B CDPO
Recall the CDPO [Mitchell, 2024] loss is given for any triplet $(x,y_{w},y_{l})$ by
@@ Appendix B CDPO
Appendix C DPO (PL)
In this setting, we are given a dataset of the form $\mathcal{D}=\{(x,y_{1},y_{2},\ldots,y_{N})\}$. In order to recover, the DPO with Plackett-Luce [Rafailov et al., 2023], we need to generalize the definition of the probability vector $p_{\theta}$ from pairs as in (7) to the following $N-1$ subsets of the list, namely the first $N$, then the first $N-1$, then first $N-2$ and so on. More precisely, for any $1\leq n<N$ we define
@@ Appendix D RPO and Distilled DPO
To recover the Distilled DPO (Equation (7) of  Fisch et al. [2024]), we consider the soft labels $p=\text{softmax}(s_{w},s_{l})$ with the loss of (8)
@@
Appendix E Proof of Lemma 5.1
First, we recall the lemma. See 5.1
@@
Appendix F C-3DPO implementation details
We show implementation details of C-3DPO-Log-$\ell_{1}$, C-3DPO-Log-$\ell_{2}$, $-\mathcal{I}$-$\ell_{1}$, and $-\mathcal{I}$-$\ell_{2}$ below.
@@ Appendix F C-3DPO implementation details
    elif algo_name == ’c3dpo_log_l2’:
@@
    elif algo_name == ’c3dpo_i_l1’:
@@
    elif algo_name == ’c3dpo_i_l2’:
@@
Appendix G Ultrafeedback Binarized Claude 3.5 Sonnet v2 win rate prompt and hyperparameters
@@
Appendix H Reddit TL;DR Claude 3.5 Sonnet v2 win rate prompt and hyperparameters
@@
Appendix I Additional win rates analysis of C-3DPO with Zephyr-7B-SFT aligned on Ultrafeedback Binarized
Figure 7: Win rates comparison of Zephyr-7B-SFT aligned on Ultrafeedback Binarized using DPO and C-3DPO. The first 4 plots show win rates of C-3DPO at individual checkpoints across 10 different inference runs. The last plot shows mean and standard error of win rates across all checkpoints and all inference runs.

### v2

@@ 7 Related Work
A few recent papers have proposed a unifying perspective on DPO-style algorithms. Notably, Tang et al. (2024) presented a generalization of DPO where different supervised learning losses are applied to the difference of implicit rewards $\big{(}\hat{r}(x,y_{w})-\hat{r}(x,y_{l})\big{)}/\beta$. In contrast, we make a clear connection between DPO-style algorithms and classification, and also extend our results to lists and auxiliary information, as do similar efforts by (Su et al., 2025; Zhao et al., 2025; Im and Li, 2024; Yao et al., 2025). A second notable example was to generalize from KL to any f-divergence when measuring the discrepancy between the target and reference model (Han et al., 2024). We also note that the first ingredient of our classification framework - $\pi_{\theta}$ was used by Sharifnassab et al. (2024) to propose a soft version of DPO.
Earlier work studied the decrease of likelihood during training (Feng et al., 2024; Xie et al., 2024), explained this phenomenon from different angles, and proposed different losses to address it. Here we provide a brief overview of a number of these results, especially those that we experimentally compare against. Pal et al. (2024) proposed DPOP which addresses the phenomenon by adding the penalty term $\max(0,-\hat{r}(x,y_{w})/\beta)$ within the log-sigmoid of the DPO loss (4). DPOP can also be viewed as DPO with a modified BT model. In this sense, it has similarities with the $\alpha$-DPO (Shen et al., 2024) (see also (Choi et al., 2025; Shao et al., 2025)).
Xiao et al. (2024) attributed the undesirable behavior to the contrastive loss of DPO not being scale-calibrated, i.e., ignoring the absolute values of implicit rewards $\hat{r}(x,y_{w})$ and $\hat{r}(x,y_{l})$. They address this by constraining the implicit rewards to a scale that matches the ground-truth reward $r$. Thus, in their proposed loss, Cal-DPO, they add the square-loss $(\hat{r}(x,y)-r(x,y))^{2},\;y\in\{y_{w},y_{l}\}$ to the DPO loss (without $\beta$). Of course, since they do not have access to the ground-truth reward, they replace it with $1/2$ and $-1/2$ for $y_{w}$ and $y_{l}$, respectively. A loss similar to Cal-DPO was proposed by Wu et al. (2024) and they named it SPPO. It is simply Cal-DPO without the DPO loss. Finally, D’Oosterlinck et al. (2024) proposed APO, which offer fine-grained control over the implicit rewards.
@@ 8 Conclusion & Future Work
In this work, we revisited the derivation of DPO and revealed two counter-intuitive findings. We first proved that the standard DPO loss can arise from a formulation that regularizes only in-sample responses. This, in turn, uncovered another surprising result: both preferred and rejected responses are likely to experience a decrease in likelihood. This insight shed light on the likelihood-displacement phenomenon observed in prior empirical studies. We then generalized this result through a unifying classification framework, highlighting a broader absence of out-of-sample KL regularization.
Building on these findings, we then introduced Constrained Controlled DPO (C2-DPO), a principled extension of DPO that controls probability displacement by constraining the redistribution of likelihood between preferred and rejected responses. C2-DPO retains the computational efficiency of standard DPO while still offering a clear RLHF interpretation.
We know that DPO displaces probabilities to unseen samples, but we do not have a clear picture in terms of the kinds of responses for which the probability increases. While Fisch et al. (2024) and Razin et al. (2025) present some findings for special cases, it would be interesting to do a more systematic study of the kinds of unseen responses whose likelihood increases during training.
Appendix A Proof of Lemmas 3.1 and 3.2
In order to prove Lemma 3.1 we will need first the following technical result. We recall that $\Delta_{n}$ denotes the simplex in $\mathbb{R}^{n}$ meaning the set of all vectors $p\in\mathbb{R}^{n}$ for which $\sum_{i=1}^{n}p_{i}=1$ and $p\geq 0$.
Lemma A.1.
Let $q\in\Delta_{n}$ and $r\in\mathbb{R}^{n}$. Let $S\subseteq[n]:=\{1,2,\ldots,n\}$ be a subset for which $|S|\geq 2$. Any optimal solution $p$ of the following problem
$\operatorname{arg\,max}_{p\in\Delta_{n}}\left\{r^{T}p-\beta\sum_{i\in S}p_{i}\ln\frac{p_{i}}{q_{i}}\right\}\;.$ (16)
satisfies that $\beta\left(\ln\left(p_{i}/q_{i}\right)-\ln\left(p_{j}/q_{j}\right)\right)=r_{i}-r_{j}$ holds true for any $i,j\in S$.
Proof.
The problem (16) is convex and Slater’s condition is satisfied; hence, the set of optimal solutions coincides with the set of KKT points. Therefore, we will consider the Lagrangian
$L(p,\lambda,\mu)=r^{T}p-\beta\sum_{i\in S}p_{i}\ln\frac{p_{i}}{q_{i}}+\lambda^{T}p+\mu\left(\sum_{i=1}^{n}p_{i}-1\right)\;,$ (17)
where $\lambda\in\mathbb{R}_{+}^{n}$ and $\mu\in\mathbb{R}$ are the Lagrange multipliers. The KKT conditions are
$\begin{cases}\frac{\partial L}{\partial p_{i}}=r_{i}-\beta\left(\ln\frac{p_{i}}{q_{i}}+1\right)+\lambda_{i}+\mu=0,&\forall i\in S\\ \frac{\partial L}{\partial p_{i}}=r_{i}+\lambda_{i}+\mu=0,&\forall i\notin S,\\ p_{i}\geq 0,\quad\sum_{i=1}^{n}p_{i}=1,&\text{(feasibility)},\\ \lambda_{i}\geq 0,&\forall i\in[n],\quad\text{(multipliers)},\\ \lambda_{i}p_{i}=0,&\forall i\in[n],\quad\text{(complementary slackness)}.\end{cases}$ (18)
Let $i\in S$, after rearranging the condition we get that $p_{i}=q_{i}e^{(r_{i}+\lambda_{i}+\mu)/\beta-1}$. Moreover, if $q_{i}>0$ then $p_{i}>0$ and hence $\lambda_{i}=0$. On the other hand, if $q_{i}=0$ then $p_{i}=0$ and the value of $\lambda_{i}$ does not matter. Hence, for simplicity, we can take $\lambda_{i}=0$ for all $i\in S$. Thus, the KKT conditions can be equivalently written as follows
$\begin{cases}p_{i}=q_{i}e^{(r_{i}+\mu)/\beta-1},&\forall i\in S.\\ r_{i}+\lambda_{i}+\mu=0,&\forall i\notin S,\\ p\in\Delta_{n},&\text{(feasibility)},\\ \lambda_{i}\geq 0,&\forall i\notin S,\quad\text{(multipliers)},\\ \lambda_{i}p_{i}=0,&\forall i\notin S,\quad\text{(complementary slackness)}.\end{cases}$ (19)
Let $\hat{r}=\max_{i\notin S}r_{i}$. We now split the proof into the two cases:
  • Case I: $\hat{r}>\beta\ln\sum_{i\in S}q_{i}e^{r_{i}/\beta-1}$
    First, for any $i\notin S$, since $\lambda_{i}\geq 0$ we have that $\mu\leq-r_{i}$ and therefore in particular $\mu\leq-\hat{r}$. Thus, there must be $i\notin S$ for which $p_{i}>0$. Indeed, if this is not the case then $p_{i}=0$ for all $i\notin S$ and we get a contradiction since (recall that $\mu\leq-\hat{r}$)
    $1=\sum_{i=1}^{n}p_{i}=\sum_{i\in S}p_{i}=\sum_{i\in S}q_{i}e^{(r_{i}+\mu)/\beta-1}\leq\sum_{i\in S}q_{i}e^{(r_{i}-\hat{r})/\beta-1}<1\ ,$ (20)
    where the last inequality follows from the condition of Case I. Therefore, for simplicity we take $i_{\ast}\notin S$ for which $p_{i_{\ast}}>0$. Therefore, $\lambda_{i_{\ast}}=0$.
    Now, for all $j\notin S$, we have $r_{i_{\ast}}+\lambda_{i_{\ast}}=r_{j}+\lambda_{j}$ and thus $r_{i_{\ast}}=r_{j}+\lambda_{j}\geq r_{j}$, which means that $r_{i_{\ast}}=\hat{r}$. Hence $\mu=-\hat{r}$, and therefore $p_{i}=q_{i}e^{(r_{i}-\hat{r})/\beta-1}$ for all $i\in S$.
    Moreover, this shows that if for some $i\notin S$ we have $r_{i}<\hat{r}$, then $p_{i}=0$. Therefore, by using the feasibility condition we obtain that any KKT point can be described by
    $p_{i}$ $=q_{i}e^{(r_{i}-\hat{r})/\beta-1},\quad\forall i\in S$
    $p_{i}$ $=0,\quad\forall i\notin S,r_{i}<\hat{r}$
    $\sum_{j\notin S,r_{j}=\hat{r}}p_{j}$ $=1-\sum_{j\in S}q_{j}e^{(r_{j}-\hat{r})/\beta-1}\;.$
  • Case II: $\hat{r}\leq\beta\ln\sum_{i\in S}q_{i}e^{r_{i}/\beta-1}$
    In this case, we will first prove that $p_{i}=0$ for all $i\notin S$. Suppose in contradiction that $p_{\ell}>0$ for some $\ell\notin S$. Then, $\lambda_{\ell}=0$ and hence $\mu=-r_{\ell}$. Recall that $r_{\ell}\leq\hat{r}$, we obtain that
    $1=\sum_{i=1}^{n}p_{i}\geq p_{\ell}+\sum_{i\in S}p_{i}=p_{\ell}+\sum_{i\in S}q_{i}e^{(r_{i}-r_{\ell})/\beta-1}\geq p_{\ell}+\sum_{i\in S}q_{i}e^{(r_{i}-\hat{r})/\beta-1}\geq p_{\ell}+1>1\;,$
    where the third inequality follows from the condition of Case II. Therefore, the unique KKT point is given by
    $p_{i}$ $=\frac{q_{i}e^{r_{i}/\beta}}{\sum_{j\in S}q_{j}e^{r_{j}/\beta}},\quad\forall i\in S$
    $p_{i}$ $=0,\quad\forall i\notin S\;.$
Finally, we see that in both cases, we have, for any $i,j\in S$, that
$\beta\left(\ln\left(p_{i}/q_{i}\right)-\ln\left(p_{j}/q_{j}\right)\right)=r_{i}-r_{j},$
as required. ∎
Now, we can prove Lemma 3.1. To this end, we recall the new optimization problem that we study
$\!\!\!\!\max_{\theta}\ \mathbb{E}_{x}\left[\mathbb{E}_{y\sim\pi_{\theta}}\big{[}r_{\phi}(x,y)\big{]}-\beta\sum_{y\in S_{x}}\pi_{\theta}(y|x)\log\frac{\pi_{\theta}(y|x)}{\pi_{\textrm{ref}}(y|x)}\right].$ (21)
Proof of Lemma 3.1. From Lemma A.1, the optimization problem (21) admits a closed-form solution
$\pi_{\theta}$, such that for any two responses $y,y^{\prime}$ within the set $S_{x}$ satisfy
$r_{\phi}(x,y)-r_{\phi}(x,y^{\prime})=\beta\ln{\frac{\pi_{\theta}(y|x)}{\pi_{\textrm{ref}}(y|x)}}-\beta\ln\frac{\pi_{\theta}(y^{\prime}|x)}{\pi_{\textrm{ref}}(y^{\prime}|x)}\;.$ (22)
Therefore, in particular, for any triplet in the preference dataset $(x,y_{w},y_{l})\in\mathcal{D}$, the BT model gives us
$p(y_{w}\succ y_{l}|x)=\sigma\left(\beta\ln{\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}}-\beta\ln\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)\;,$ (23)
which are exactly the probabilities used in the DPO derivation, and taking the negative-log-likelihood of (23) over the entire dataset yields exactly the DPO loss. ∎
We complete this section with the proof of Lemma 3.2.
Proof of Lemma 3.2. If $r_{\phi}(x,y)\leq\hat{r}_{x}$ for every $y\in S_{x}$, than the condition $\hat{r}_{x}>\beta\ln(\sum_{y\in S_{x}}\pi_{\textrm{ref}}(y|x)e^{r_{\phi}(x,y)/\beta-1})$ is satisfied. This is because
$\beta\ln\sum_{y\in S_{x}}\pi_{\textrm{ref}}(y|x)e^{r_{\phi}(x,y)/\beta-1}$ $=\beta\ln\sum_{y\in S_{x}}\pi_{\textrm{ref}}(y|x)e^{\hat{r}_{x}/\beta+(r_{\phi}(x,y)-\hat{r}_{x})/\beta-1}$
$=\hat{r}_{x}+\beta\ln\sum_{y\in S_{x}}\pi_{\textrm{ref}}(y|x)e^{(r_{\phi}(x,y)-\hat{r}_{x})/\beta-1}$
$\leq\hat{r}_{x}+\beta\ln\sum_{y\in S_{x}}\pi_{\textrm{ref}}(y|x)e^{-1}$
$=\hat{r}_{x}+\beta\left(-1+\ln\sum_{y\in S_{x}}\pi_{\textrm{ref}}(y|x)\right)$
$\leq\hat{r}_{x}-\beta\ ,$
where the last inequality follows from the fact that $\sum_{y\in S_{x}}\pi_{\textrm{ref}}(y|x)\leq 1$. This proves that the condition holds true.
Therefore, as we saw in the proof of Lemma 3.1, in any optimal solution $\pi_{\theta}$ we have for all $y\in S_{x}$, that
$\pi_{\theta}(y|x)=\pi_{\textrm{ref}}(y|x)e^{(r_{\phi}(x,y)-\hat{r}_{x})/\beta-1}\leq\pi_{\textrm{ref}}(y|x)e^{-1}=e^{-1}\pi_{\textrm{ref}}(y|x),$
which proves the desired result. ∎
Appendix B Derivations of Existing Algorithms via the Classification Framework
B.1 IPO
@@
B.2 CDPO
Recall the CDPO Mitchell [2024] loss is given for any triplet $(x,y_{w},y_{l})$ by
@@ B.2 CDPO
B.3 DPO (PL)
In this setting, we are given a dataset of the form $\mathcal{D}=\{(x,y_{1},y_{2},\ldots,y_{N})\}$. In order to recover, the DPO with Plackett-Luce Rafailov et al. [2023], we need to generalize the definition of the probability vector $p_{\theta}$ from pairs as in (7) to the following $N-1$ subsets of the list, namely the first $N$, then the first $N-1$, then first $N-2$ and so on. More precisely, for any $1\leq n<N$ we define
@@ B.4 RPO and Distilled DPO
To recover the Distilled DPO (Equation (7) of  Fisch et al. [2024]), we consider the soft labels $p=\text{softmax}(s_{w},s_{l})$ with the loss of IPO.
@@
B.5 List of Preferences
Our classification framework can be extended to work with lists, rather than pairs, of preferences. In particular, assume that we have $N$ responses for each prompt $x$, giving us a dataset of the form $\mathcal{D}=\{(x,y_{1},y_{2},\ldots,y_{N})\}.$ In this case, we can define a list version of the probability vector $p_{\theta}(x,y_{1},\ldots,y_{N})$ similarly to (6), together with a target distribution $p$. With this simple extension we can now incorporate existing DPO-style algorithms that work with lists. For instance, we can show that DPO with the Plackett-Luce model for preferences Rafailov et al. [2023] can be captured in our classification framework, again using hard labels and the CE loss. See Appendix B.3 for a proof.
B.6 Auxiliary Information
More formally, suppose now that our dataset is comprised of ${\mathcal{D}}=(x,y_{w},s_{w},y_{l},s_{l})$. To obtain the soft labels we can employ, for instance, $p=\text{softmax}(s_{w},s_{l})$. Combining this with the IPO loss, we recover Distilled DPO (Eq. 7 in Fisch et al. 2024). Using the same soft labels, but with the CE loss recovers RPO (see Sec. 3.3.2 in Adler et al. 2024). We provide more details on both algorithms in Appendix B.4. These natural extensions further demonstrate that our classification framework is fairly general as well as sufficiently flexible to capture a large number of existing DPO-style algorithms.
Appendix C A Concrete Example Showing that the Original DPO-style Problem is Under-constrained
Below we give a concrete example further highlighting that the winner-loser probabilities can move in arbitrary directions.
Remark C.1 (A Concrete Example).
Suppose we have $\varepsilon=1/11$ and for simplicity we set $\beta=1$, $\pi_{\textrm{ref}}(y_{w}|x)=0.02$, and $\pi_{\textrm{ref}}(y_{l}|x)=0.01$, which result in $\eta=20$. (See Eqn 8 for the definition of $\eta$) We identify two pairs of probabilities that satisfy (8):
  1. 1.
[…]
  2. 2.
[…]
Appendix D Proofs of Section 5
D.1 Proof of Proposition 5.1
See 5.1
Proof.
For simplicity we assume that $\varphi$ is monotonically increasing (the same arguments can be easily applied to the monotonically decreasing case). Since $\varepsilon<1/2$, we have that $(1-\varepsilon)/\varepsilon>1$, and therefore from (10), we get that
We will show that $\pi_{\theta^{*}}(y_{l}|x)<\pi_{\textrm{ref}}(y_{l}|x)$ by contradiction. To this end, we assume that $\pi_{\theta^{*}}(y_{l}|x)\geq\pi_{\textrm{ref}}(y_{l}|x)$. From (26), we immediately obtain that $\pi_{\theta^{*}}(y_{w}|x)>\pi_{\textrm{ref}}(y_{w}|x)$. Applying the constraint function $\varphi$ to the following two inequalities:
and using its monotonicity, we obtain
$\varphi\big{(}\pi_{\theta^{*}}(y_{w}|x)\big{)}>\varphi\big{(}\pi_{\textrm{ref}}(y_{w}|x)\big{)}\quad\text{and}\quad\varphi\big{(}\pi_{\theta^{*}}(y_{l}|x)\big{)}\geq\varphi\big{(}\pi_{\textrm{ref}}(y_{l}|x)\big{)}.$
By adding both sides of the above two inequalities, we get a contradiction to (9). Thus, proving that $\pi_{\theta^{*}}(y_{l}|x)<\pi_{\textrm{ref}}(y_{l}|x)$. Similarly, we can prove that $\pi_{\theta^{*}}(y_{w}|x)>\pi_{\textrm{ref}}(y_{w}|x)$. ∎
D.2 Proof of Lemma 5.1
See 5.1
@@ Appendix D Proofs of Section 5
Appendix E C2-DPO Implementation Details
We show implementation details of C2-DPO-Log-$\ell_{1}$, C2-DPO-Log-$\ell_{2}$, $-\mathcal{I}$-$\ell_{1}$, and $-\mathcal{I}$-$\ell_{2}$ below.
@@ Appendix E C2-DPO Implementation Details
    elif algo_name == ’c2dpo_log_l2’:
@@
    elif algo_name == ’c2dpo_i_l1’:
@@
    elif algo_name == ’c2dpo_i_l2’:
@@
Appendix F Ultrafeedback Binarized Claude 3.5 Sonnet v2 win rate prompt and hyperparameters
@@
Appendix G Reddit TL;DR Claude 3.5 Sonnet v2 win rate prompt and hyperparameters
@@
Figure 7:
Appendix H Additional win rates analysis of C-3DPO with Zephyr-7B-SFT aligned on Ultrafeedback Binarized
Figure 8: Win rates comparison of Zephyr-7B-SFT aligned on Ultrafeedback Binarized using DPO and C-3DPO. The first 4 plots show win rates of C-2DPO at individual checkpoints across 10 different inference runs. The last plot shows mean and standard error of win rates across all checkpoints and all inference runs.